OpenStudy (anonymous):

WHAT IS The range of 3 √ x !?!?! PLEASE PLEASE PLAESE PLAESE I BEGGG I BEGGGGGGG! LOL

OpenStudy (kainui):

What can x be? What numbers can you not take the square root of?

OpenStudy (anonymous):

O?

OpenStudy (anonymous):

\[\sqrt{0} = 0\]

OpenStudy (kainui):

You can take the square root of 0, since 0*0=0. You can't divide by 0, don't forget that when determining range and domain.

OpenStudy (anonymous):

YES BUT LET ME SHOW U GUYS THE OPTIONS

OpenStudy (kainui):

Try turning off your caps lock, it doesn't really help you.

OpenStudy (anonymous):

\[\sqrt{-1} \notin R\]

OpenStudy (anonymous):

o sorry lol :)

OpenStudy (anonymous):

what types of numbers can we not square root?

OpenStudy (kainui):

Haha it's fine. So yes, eigenschmeigen is right, if you can guess what that means. Remember that a domain is all the numbers of x you can choose to satisfy an equation while the range is all the numbers you can have as a result of your domain.

OpenStudy (anonymous):

i attached a picture of the options can u guys check it out

OpenStudy (kainui):

Numbers in these brackets [x,y] mean that they are included in the range or domain. Numbers in parenthesis (x,y) mean that they're not included, but every number between them is.

OpenStudy (anonymous):

can u guys help because he didnt help :(

OpenStudy (anonymous):

Well, what do you think the answer is?

OpenStudy (anonymous):

@Kainui i personally think you were explaining very well

OpenStudy (anonymous):

C ?

OpenStudy (anonymous):

what happens if you put a negative number for x?

OpenStudy (anonymous):

@eigenschmeigen im just very slow in math like i dont get anything im sorry and i need to get this half a credit graduate with my class by june 3rd and its a whole bunch of quizes this invluded and im just stressing! :(((

OpenStudy (anonymous):

what happens if we put a negative number in for x?

OpenStudy (anonymous):

what do we get, for example, if i put in -16

OpenStudy (anonymous):

i believe the answer is A.

OpenStudy (anonymous):

A....

OpenStudy (anonymous):

yes you would be correct

OpenStudy (anonymous):

thanks so much.

OpenStudy (anonymous):

I hope you're not just guessing and you truly, genuinely understand the concept behind your answer.